OpenStudy (auw7):

simplify (x^3/5)^2/3 I got x^6/15, but I don't think its that easy... I will award a medal!

OpenStudy (welshfella):

you multiply the exponents

OpenStudy (welshfella):

yu are correct

OpenStudy (auw7):

oh thank you!

OpenStudy (anonymous):

It is that easy. You just need to reduce

OpenStudy (welshfella):

3/5 * 2/3 = 6/15

OpenStudy (auw7):

okay thanks! I did that but I thought that maybe I was doing it wrong. Thank you

OpenStudy (welshfella):

you can reduce it ti simplest form 6 and 15 are both divisible by 3

OpenStudy (anonymous):

\[\frac{ 3 }{ 5 }\times \frac{ 2 }{ 3 }=\frac{ 2 }{ 5 }\]